Keith’s Florists has 13 delivery trucks, used mainly to deliver flowers and flower arrangements in the Greenville, South Carolina, area. Of these 13 trucks, 6 have brake problems. A sample of 5 trucks is randomly selected. What is the probability that 2 of those tested have defective brakes?
Answer #1


Here, n = 5, p = 0.4615, (1 - p) = 0.5385 and x = 2
As per binomial distribution formula P(X = x) = nCx * p^x * (1 - p)^(n - x)

We need to calculate P(X = 2)
P(X = 2) = 5C2 * 0.4615^2 * 0.5385^3
P(X = 2) = 0.3326
0